OpenStudy (anonymous):

Solve 7.74^x + 3 = 10.2

OpenStudy (imstuck):

Solving for x? And I bet that you're studying natural logs, right?

OpenStudy (imstuck):

common logs, is more like it!

OpenStudy (imstuck):

take the log of each side, like this:

OpenStudy (imstuck):

\[\log 7.74^{x}=\log(10.2-3)\]

OpenStudy (imstuck):

\[\log 7.74^{x}=\log(7.2)\]Log laws tell us that we can move that x exponent to the front to obtain this:

OpenStudy (imstuck):

\[x \log (7.74)=\log (7.2)\]in order to solve for x, we will divide log(7.2) by log(7,74) like this:

OpenStudy (anonymous):

ok

OpenStudy (imstuck):

\[x=\frac{ \log (7.2) }{ \log (7.74) }\]\[x=\frac{ .857332 }{ .888741 }\]

OpenStudy (imstuck):

What do you get when you divide that out?

OpenStudy (imstuck):

you will find the values for those logs on your calculator.

OpenStudy (anonymous):

.9646595964

OpenStudy (imstuck):

Yep! Although you probably don't need it to all those decimal places! Talk about accuracy!

OpenStudy (anonymous):

x ≈ -2.15 x ≈ -0.47 x ≈ 0.03 x ≈ 0.28

OpenStudy (imstuck):

eek, let me look at that again, then!

OpenStudy (anonymous):

okie dokie

OpenStudy (imstuck):

hmmm...hold on for a bit while I review. I cannot see any other way to solve this for x.

OpenStudy (anonymous):

ok

OpenStudy (imstuck):

Every place i look tells me the same thing...to take the common log of both sides. Let's ask @D3xt3R . He's really good with this. Maybe he can lend us his brain to get another way.

OpenStudy (imstuck):

Logs exist because of these types of problems

OpenStudy (anonymous):

;D ... Hello guys ;)

OpenStudy (anonymous):

uh oh Solve 7.7^4x + 3 = 10.2

OpenStudy (anonymous):

i typed the problem worng

OpenStudy (imstuck):

help us to see the error in our ways, @D3xt3R !

OpenStudy (imstuck):

oh my! You typed the problem wrong! I'm so glad to hear that cuz I thought I had gone stupid, or something! :/

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

Is\[7.7^{4x}+3=10.2~~or~~7.74^x+3=10.2???\]

OpenStudy (imstuck):

That is so much better! Now the problem is this after we take the common log of both sides:\[4x \log (7.74)=\log (7.2)\]

OpenStudy (imstuck):

4x = .96466

OpenStudy (imstuck):

Now solve for x.

OpenStudy (imstuck):

x = .24116

OpenStudy (imstuck):

hmm...still off some

OpenStudy (anonymous):

No matter what log do you chose, the result will be the same

OpenStudy (anonymous):

\[ln(7.7^{4x})=ln(7.2)\]\[4x=\frac{ln(7.2)}{ln(7.7)}\]\[x=0.24178\]

OpenStudy (imstuck):

see? it's not coming out right still

OpenStudy (anonymous):

≈ -2.15 x ≈ -0.47 x ≈ 0.03 x ≈ 0.28

OpenStudy (anonymous):

yeah??

OpenStudy (anonymous):

Thinking

OpenStudy (anonymous):

I got it ;)

OpenStudy (anonymous):

whattt

OpenStudy (anonymous):

\[7.7^{4x+3}=10.2\]\[ln(7.7^{4x+3})=ln(10.2)\]\[4x+3=\frac{ln(10.2)}{ln(7.7)}\]\[4x=1.134864-3\]\[4x=-1.865136\]\[x=-\frac{1.865136}{4}\]@IMStuck and @jonnyrco The correct answer is:\[\boxed{x\approx-0.47}\]

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

thanks

OpenStudy (imstuck):

aha those darn old exponents!

OpenStudy (anonymous):

haha